# From Estimands to Robust Inference of Treatment Effects in Platform Trials

### v1

@@ 4.3 Post-stratification
An alternative method to account for the varying $\pi_{j}({Z})$ is to stratify subjects based on the values that $\pi_{j}({Z})$ can take. This approach is known as post-stratification in sample surveys (Fuller,, 2009) or sub-classification in the causal inference literature (Rosenbaum and Rubin,, 1984). Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible patients ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$. The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
@@
where $n({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$. This estimator does not need stabilizing, as the sum of weights is
$\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})=1.$
Compared with the SIPW, the PS requires an extra step of constructing strata for every pair $j$ and $k$.
The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time. For instance, in Figure 1, ${Z}$ includes both the disease subtype and the enrollment window; in the SIMPLIFY example in Figure 2, $Z$ is the indicator of HS and/or DA and there is only a single enrollment window. On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary. Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$. It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section 7.1.
The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$. To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages. This gives the following adjusted post-stratification (APS) estimator of $\theta_{jk}$,
@@
Asymptotically, the PS and APS estimators are shown to be equivalent to certain types of SAIPW estimators, as detailed in Section 5.
4.4 Summary of estimators
[…]
@@ 5.1 Asymptotic normality
For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$. If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in S1 of the Supplementary Material.
Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions. In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
@@
Under Assumption 1, and Assumption 2 if adjusting for covariates, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$ whose explicit expression is given in each of the following specific case.
@@
where $S$ is the covariate for post-stratification, i.e., for patient $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplementary Material.
@@
All technical proofs are given in the Supplementary Material. The results in Theorem 1 for estimators with covariate adjustment hold regardless of whether the working model is correct or not. The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators in Section 6.
@@ 5.2 Efficiency comparison
As the six estimators proposed in Section 4 have different asymptotic covariance matrices, we compare the asymptotic relative efficiency between pairs of these estimators in the subsequent corollaries. The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
@@
Under Assumption 1, the difference in asymptotic covariance matrices of the IPW and SIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.$
Since the matrix $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ in Corolary 3 is positive semidefinite (shown in S6 of the Supplementary Material) but ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ may have arbitrary signs, the asymptotic relative efficiency of IPW versus SIPW is indefinite. When ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ are both non-negative, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ is positive semidefinite and, hence, SIPW always has equal or better asymptotic efficiency than IPW in estimating any differentiable function of $\bm{\vartheta}_{jk}$. On the other hand, an example of estimating $\theta_{jk}-\theta_{kj}$ is shown in S6 of the Supplement Material in which IPW is asymptotically more efficient than SIPW.
Nonetheless, SIPW is often more efficient in simulations and real data applications (see Section 7). SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change the SIPW estimator $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect the IPW analogue $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023). For these reasons, we do not consider IPW in the remaining efficiency comparisons.
@@
Under Assumptions 1-2, the difference in asymptotic covariance matrices of the AIPW and SAIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.\vspace{-2mm}$
Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition
$\delta_{jk}=\delta_{kj}=0,\vspace{-3mm}\vspace{-2mm}$ (4)
which is known as the prediction unbiasedness condition (Guo and Basse,, 2021) at the population level, where $\delta_{jk}=E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\}$ and $\delta_{kj}=E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\}$. Condition (4) holds when the working models used for covariate adjustment are correctly specified. It also holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified. Later discussions about adjustment (8) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet condition (4), it can be readily modified by (8) to satisfy this requirement. Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
The next two results show when covariate adjustment leads to an efficiency gain. Corollary 3 compares the two stabilized IPW estimators, with or without covariate adjustment.
@@
When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect. Conditions (5)-(7) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the platform trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent. Actually, condition (5) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
@@
where $n_{jk}(S)$ is the number of elements in $\{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S\}$. The corresponding limit of $\tilde{\mu}_{jk}(\bm{X})$ in Assumption 2 is $\mu_{jk}(\bm{X})+E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}$ satisfying (5) with $\delta_{jk}=\delta_{kj}=0$. Conditions (6)-(7) hold when working models are correct. Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (6)-(7) embedded. For example, (5)-(7) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2023); see S10 in the Supplement Material for details.
@@
Corollary 4 compares the two post-stratification estimators, with or without additional adjustment for covariates.
@@
Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in the trivial scenario where either no covariate adjustment beyond $S$ in one of the treatment arms or $\pi_{j}(S)+\pi_{k}(S)=1$ and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent conditioned on $S$.
To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting. For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$. This leads to the following result.
@@
Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,ps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$ when $\bm{X}=S$.
Corollary 5 states that the PS estimator that stratifies on $S$ is asymptotically equivalent to the SAIPW estimator that adjusts for $S$ (as well as the corresponding AIPW estimator due to its equivalence with the SAIPW when (4) holds).
The next corollary compares APS and SAIPW with the same covariate adjustment. It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (5), the APS estimator is asymptotically equivalent to the SAIPW estimator.
@@
Under Assumptions 1-2 and condition (5), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,aps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$.
The last corollary completes the comparison around SAIPW with $\bm{X}=S$ and shows that under (5)-(7), adjusting for more covariates beyond just $S$ leads to additional asymptotic efficiency gain.
Corollary 7 (Comparison of SAIPW and SAIPW with $\bm{X}=S$).
Under Assumptions 1-2 and conditions (5)-(7), SAIPW with $S\subset\bm{X}$ is more efficient than SAIPW with $\bm{X}=S$.

### v4

@@ 4.4 Post-stratification
When $\pi_{j}({Z})$ takes discrete values, an alternative method to account for the varying $\pi_{j}({Z})$ is to stratify individuals based on the values that $\pi_{j}({Z})$ can take. This approach is known as post-stratification in sample surveys (Fuller,, 2009). Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible individuals in ${{{\cal I}}}_{jk}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$. The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
@@
where $n({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$. This estimator does not need stabilizing, as the sum of weights in $\hat{\theta}_{jk}^{\rm\,ps}$ is equal to 1.
The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time. On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary. Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$. It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section S2 of the Supplement.
The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$. To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages. This gives the adjusted post-stratification (APS) estimator of $\theta_{jk}$,
@@ 5.1 Asymptotic normality
For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$. If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in Section S1 of the Supplement.
Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions. In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$. Similarly, $E_{jk}(\cdot\mid S)=E_{kj}(\cdot\mid S)$, ${\rm Var}_{jk}(\cdot\mid S)={\rm Var}_{kj}(\cdot\mid S)$, and ${\rm Cov}_{jk}(\cdot\mid S)={\rm Cov}_{kj}(\cdot\mid S)$ denote these quantities further conditioned on $S$, where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$.
@@
Under Assumption 1, and Assumption 2 if covariates are adjusted, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$, where $\hat{\bm{\vartheta}}_{jk}^{\star}=(\hat{\theta}_{jk}^{\star},\hat{\theta}_{kj}^{\star})^{T}$, $\bm{a}^{T}$ is the vector transpose of $\bm{a}$, and the explicit expression of $\bm{\Sigma}^{\star}_{jk}$ is given in each of the following specific case.
@@
where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplement.
@@
All technical proofs are given in the Supplement. The results in Theorem 1 and the efficiency comparison in Section 5.2 are new to the literature and differ from existing results for IPW or AIPW estimators using alternative weights, such as estimated propensity scores (Robins et al.,, 1994; Lunceford and Davidian,, 2004). We also show that the AIPW and SAIPW estimators with $\mu_{jk}=E_{jk}(Y^{(j)}\mid\bm{X})$ and $\mu_{kj}=E_{jk}(Y^{(k)}\mid\bm{X})$ are asymptotically equivalent and both attain semiparametric efficiency. The proof is provided in Section S3 of the Supplement. Furthermore, our results include PS and APS estimators and hold regardless of whether the working model is correctly specified. The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators presented in Section S1 of the Supplement.
@@ 5.2 Efficiency comparison
We compare the asymptotic relative efficiency between pairs of six estimators in the subsequent corollaries. The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
@@
Under Assumption 1, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
Although $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ is positive semidefinite, the asymptotic relative efficiency of IPW versus SIPW is indefinite because the covariances may have arbitrary signs; see Section S3 of the Supplement for more discussions. Nonetheless, SIPW is often more efficient in simulations (see Section S2 of the Supplement). SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023). For these reasons, we do not consider IPW in the remaining efficiency comparisons.
@@
Under Assumptions 1-2, $\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},}$ ${\textstyle{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition $\delta_{jk}=\delta_{kj}=0$, which holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified. Later discussions about adjustment (7) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet this condition, it can be readily modified by (7) to satisfy this requirement. Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
The next two results show when covariate adjustment leads to an efficiency gain.
@@
When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect. Conditions (4)-(6) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the master protocol trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent. Condition (4) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
@@
where $n_{jk}(S)$ is the number of elements in $\{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S\}$. The corresponding limit of $\tilde{\mu}_{jk}(\bm{X})$ in Assumption 2 is $\mu_{jk}(\bm{X})+E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}$ satisfying (4) with $\delta_{jk}=\delta_{kj}=0$. Conditions (5)-(6) hold when working models are correct. Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (5)-(6) embedded. For example, (4)-(6) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2025); see S3 in the Supplement for details.
@@
Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in some trivial scenarios.
@@
To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting. For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$. The following result shows the asymptotic equivalence between these two estimators.
@@
Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix.
The last corollary compares APS and SAIPW with the same covariate adjustment. It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (4), the APS estimator is asymptotically equivalent to the SAIPW estimator.
@@
Under Assumptions 1-2 and condition (4), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix.
It follows from Corollaries 4-6 that, under (4)-(6), SAIPW with $\bm{X}\supset S$ is more efficient than SAIPW with $\bm{X}=S$; see Section S3 in the Supplement for a proof.